# Bayesian Optimization of Robustness Measures Using Randomized GP-UCB-based Algorithms under Input Uncertainty

### v1

@@ C.5 Equality between $q(a)$ in Table 4 and $q_{1}(a)$, $q_{2}(a)$, and $q_{3}(a)$
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\frac{1}{\alpha}\int_{0}^{\alpha}(v_{f}({\bm{x}};\alpha^{\prime})-v_{f^{\prime}}({\bm{x}}^{\prime};\alpha^{\prime})){\rm d}\alpha^{\prime}$ $\leq\frac{1}{\alpha}\int_{0}^{\alpha}\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|{\rm d}\alpha^{\prime}$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]|]-\mathbb{E}_{{\bm{w}}}[|f^{\prime}({\bm{x}}^{\prime},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]|]$
@@
Let $P$ be a candidate distribution of ${\bm{w}}$, $\mathcal{A}$ be a family of candidate distributions, and $\rho_{P}(\cdot)$ be a robustness measure with respect to $P$. Suppose that $q(\cdot)$ is a non-decreasing concave function satisfying $q(0)=0$ and $|\rho_{P}(f({\bm{x}},{\bm{w}}))-\rho_{P}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|)$ for any $P$. For any $\epsilon>0$, from the definition of infimum, there exists a distribution $P^{\prime}\in\mathcal{A}$ such that
$\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\inf_{P\in\mathcal{A}}\rho_{P}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\geq\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))-\epsilon.$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $\leq\inf_{P\in\mathcal{A}}\rho_{P}(f({\bm{x}},{\bm{w}}))-\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\epsilon$
$\leq\rho_{P^{\prime}}(f({\bm{x}},{\bm{w}}))-\rho_{P^{\prime}}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\epsilon$
$\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right)+\epsilon.$
@@
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
@@
Let $\mathcal{M}$ be a $K$-Lipschitz function. Suppose that $q(a)$ is a non-decreasing concave function satisfying $q(0)=0$ and
$|\tilde{\rho}(f({\bm{x}},{\bm{w}}))-\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right),$
where $\tilde{\rho}(\cdot)$ is a robustness measure. Then, for the robustness measure $\rho(\cdot)=(\mathcal{M}\circ\tilde{\rho})(\cdot)$, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))=\mathcal{M}(\tilde{\rho}(f({\bm{x}},{\bm{w}})))-\mathcal{M}(\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$ $\leq K|\tilde{\rho}(f({\bm{x}},{\bm{w}}))-\tilde{\rho}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|$
$\leq Kq\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
@@
Let $\alpha_{1},\alpha_{2}\geq 0$. Suppose that $q_{i}(a)$ is a non-decreasing concave function satisfying $q_{i}(0)=0$ and
$|{\rho}_{i}(f({\bm{x}},{\bm{w}}))-{\rho}_{i}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))|\leq q_{i}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right),$
where $\rho_{i}(\cdot)$ is a robustness measure. Then, for the robustness measure $\rho(\cdot)=\alpha_{1}\rho_{1}(\cdot)+\alpha_{2}\rho_{2}(\cdot)$, the following holds:
$\rho(f({\bm{x}},{\bm{w}}))-\rho(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))$ $=\alpha_{1}\rho_{1}(f({\bm{x}},{\bm{w}}))+\alpha_{2}\rho_{2}(f({\bm{x}},{\bm{w}}))-(\alpha_{1}\rho_{1}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}}))+\alpha_{2}\rho_{2}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$
$=\alpha_{1}(\rho_{1}(f({\bm{x}},{\bm{w}}))-\rho_{1}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))+\alpha_{2}(\rho_{2}(f({\bm{x}},{\bm{w}}))-\rho_{2}(f^{\prime}({\bm{x}}^{\prime},{\bm{w}})))$
@@
Hence, for the function $f({\bm{x}},{\bm{w}})$ with respect to ${\bm{w}}$, each element ${\bm{x}}\in\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\}$ satisfies $f({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$. Moreover, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following inequality holds:
@@
where the third inequality is derived by $\mu_{t-1}({\bm{x}},{\bm{w}})\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$ and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho(\mu_{t-1}({\bm{x}},{\bm{w}}))=\rho(\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}}))$. Furthermore, from Assumption 4.1, there exists a function
@@
$F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})=\rho(f({\bm{x}}^{\ast},{\bm{w}}))-\rho(f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))|\right).$
@@ C.7 Upper Bound of $\mathbb{E}[L_{max}]$
Thus, for the function $f({\bm{x}},[{\bm{w}}]_{t})$ with respect to ${\bm{w}}$, $f({\bm{x}},[{\bm{w}}]_{t})\in G^{\dagger}_{t-1}({\bm{x}})$ holds. Here, from the theorem’s assumption, since ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ satisfy (A.1), the following holds:
@@
$F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})=\rho(f({\bm{x}}^{\ast},{\bm{w}}))-\rho(f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))|\right).$
@@
$=\mathbb{E}_{{\bm{w}}}[\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}})]=\rho(\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}})).$
@@
By combining these, and using the same argument as in the proof of Theorem 4.1, we obtain Theorem B.1. Theorem B.2B.4 can also be obtained by using the same argument. ∎
C.12 Proof of Theorem B.5B.8
@@
In order to show Theorem B.5B.8, we evaluate $\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]$ by using the same argument as in the proof of Theorem B.1B.4. Under the given
@@ D Experimental Details and Additional Experiments
In this section, we give details of the numerical experiments and the results of additional numerical experiments.
@@ D.1 Experimental Details
BPT-UCB (fix)
In BPT-UCB (fix), the definition of ${\rm BPTUCB}({\bm{x}})$ in BPT-UCB was changed to ${\rm BPTUCB}({\bm{x}})=\hat{p}_{t}({\bm{x}})+3\gamma_{t}({\bm{x}})$, and ${\bm{x}}_{t+1}$ and ${\bm{w}}_{t+1}$ were selected using the same procedure as BPT-UCB.
@@
BBBMOBO (fix)
In BBBMOBO (fix), we used $\beta_{t}=9$ and ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, where $\tilde{\bm{x}}_{t}$ is given in Definition 3.1.
@@
Proposed (fix)
In Proposed (fix), ${\bm{x}}_{t}$ was selected by Definition 3.1, where we used $\beta_{t}=9$.
Values of ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ in EXP, PTR and EXP-MAE were calculated by using results in Table 3. Here, ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for $-\alpha\mathbb{E}[|f({\bm{x}},{\bm{w}})-F_{{\rm exp}}({\bm{x}})|]$ were calculated by using the results of the mean absolute deviation and monotonic Lipschitz map in Table 3. Combining these and the results for the weighted sum, we calculated ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for EXP-MAE.
D.2 Additional Experiments
In this section, we changed the experiment in Section 5.1 to the uncontrollable setting, where ${\bm{w}}_{t}$ cannot be selected and is obtained randomly according to the probability mass function $p({\bm{w}})$. The method for selecting ${\bm{x}}_{t}$ was the same as in Section 5.1. As in Section 5.1, Figure 4 shows that the proposed method outperforms the comparison methods in most settings. On the other hand, in the experiments in Section 5.1, the regret of Proposed (fix) did not decrease in the 4D synthetic function (EXP), while in Figure 4, the regret decreased more than Figure 2. This is because ${\bm{w}}$ could not be selected, and as a result of random sampling, more exploration was performed, making it possible to avoid local solutions.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 4: Average regrets for 100 simulations under the uncontrollable setting using each method. The top, middle, and bottom rows show the results for the 2D, 4D, and 6D synthetic function settings, respectively, and the left, middle, and right columns show the results for EXP, PTR, and EXP-MAE, respectively. The length of the error bar represents twice the standard error.

### v2

@@ C Proofs
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})=\frac{1}{\alpha}\int_{0}^{\alpha}(v_{f}({\bm{x}};\alpha^{\prime})-v_{f^{\prime}}({\bm{x}}^{\prime};\alpha^{\prime})){\rm d}\alpha^{\prime}$ $\leq\frac{1}{\alpha}\int_{0}^{\alpha}\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|{\rm d}\alpha^{\prime}$
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})$ $=\mathbb{E}_{{\bm{w}}}[|f({\bm{x}},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f({\bm{x}},{\bm{w}})]|]-\mathbb{E}_{{\bm{w}}}[|f^{\prime}({\bm{x}}^{\prime},{\bm{w}})-\mathbb{E}_{{\bm{w}}}[f^{\prime}({\bm{x}}^{\prime},{\bm{w}})]|]$
@@
Let $P$ be a candidate distribution of ${\bm{w}}$, $\mathcal{A}$ be a family of candidate distributions, and $\rho_{P}(\cdot)$ be a robustness measure with respect to $P$. The distributionally robust measure is defined as $\rho({f({\bm{x}},\cdot)})=\inf_{P\in\mathcal{A}}\rho_{P}({f({\bm{x}},\cdot)})$. Let $q_{\mathcal{A}}(\cdot)$ be a function satisfying $|\rho_{P}({f({\bm{x}},\cdot)})-\rho_{P}({f^{\prime}({\bm{x}}^{\prime},\cdot)})|\leq q_{\mathcal{A}}(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|)$ for any $P$. Additionally, assume that $q_{\mathcal{A}}(\cdot)$ is a non-decreasing concave function satisfying $q_{\mathcal{A}}(0)=0$. For any $\epsilon>0$, from the definition of infimum, there exists a distribution $P^{\prime}\in\mathcal{A}$ such that
$\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})=\inf_{P\in\mathcal{A}}\rho_{P}({f^{\prime}({\bm{x}}^{\prime},\cdot)})\geq\rho_{P^{\prime}}({f^{\prime}({\bm{x}}^{\prime},\cdot)})-\epsilon.$
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})$ $\leq\inf_{P\in\mathcal{A}}\rho_{P}({f({\bm{x}},\cdot)})-\rho_{P^{\prime}}({f^{\prime}({\bm{x}}^{\prime},\cdot)})+\epsilon$
$\leq\rho_{P^{\prime}}({f({\bm{x}},\cdot)})-\rho_{P^{\prime}}({f^{\prime}({\bm{x}}^{\prime},\cdot)})+\epsilon$
$\leq q_{\mathcal{A}}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right)+\epsilon.$
@@
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})\leq q_{\mathcal{A}}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
This implies that $q(a)=q_{\mathcal{A}}(a)$.
@@
Let $\mathcal{M}$ be a $K$-Lipschitz function, and $\tilde{\rho}(\cdot)$ be a robustness measure. The target robustness measure is defined as $\rho(\cdot)=(\mathcal{M}\circ\tilde{\rho})(\cdot)$. Let $\tilde{q}(a)$ be a function satisfying
$|\tilde{\rho}({f({\bm{x}},\cdot)})-\tilde{\rho}({f^{\prime}({\bm{x}}^{\prime},\cdot)})|\leq\tilde{q}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Additionally, assume that $\tilde{q}(a)$ is a non-decreasing concave function satisfying $\tilde{q}(0)=0$. Then, for the robustness measure $\rho(\cdot)$, the following holds:
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})=\mathcal{M}(\tilde{\rho}({f({\bm{x}},\cdot)}))-\mathcal{M}(\tilde{\rho}({f^{\prime}({\bm{x}}^{\prime},\cdot)}))$ $\leq K|\tilde{\rho}({f({\bm{x}},\cdot)})-\tilde{\rho}({f^{\prime}({\bm{x}}^{\prime},\cdot)})|$
$\leq K\tilde{q}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
This implies that $q(a)=K\tilde{q}(a)$.
@@
Let $\alpha_{1},\alpha_{2}\geq 0$, and let $\rho_{i}(\cdot)$ be a robustness measure. The target robustness measure is defined as $\rho(\cdot)=\alpha_{1}\rho_{1}(\cdot)+\alpha_{2}\rho_{2}(\cdot)$. Let $q_{i}(a)$ be a function satisfying
$|{\rho}_{i}({f({\bm{x}},\cdot)})-{\rho}_{i}({f^{\prime}({\bm{x}}^{\prime},\cdot)})|\leq q_{i}\left(\sup_{{\bm{w}}\in\Omega}|f({\bm{x}},{\bm{w}})-f^{\prime}({\bm{x}}^{\prime},{\bm{w}})|\right).$
Additionally, assume that $q_{i}(a)$ is a non-decreasing concave function satisfying $q_{i}(0)=0$. Then, for the robustness measure $\rho(\cdot)$, the following holds:
$\rho({f({\bm{x}},\cdot)})-\rho({f^{\prime}({\bm{x}}^{\prime},\cdot)})$ $=\alpha_{1}\rho_{1}({f({\bm{x}},\cdot)})+\alpha_{2}\rho_{2}({f({\bm{x}},\cdot)})-(\alpha_{1}\rho_{1}({f^{\prime}({\bm{x}}^{\prime},\cdot)})+\alpha_{2}\rho_{2}({f^{\prime}({\bm{x}}^{\prime},\cdot)}))$
$=\alpha_{1}(\rho_{1}({f({\bm{x}},\cdot)})-\rho_{1}({f^{\prime}({\bm{x}}^{\prime},\cdot)}))+\alpha_{2}(\rho_{2}({f({\bm{x}},\cdot)})-\rho_{2}({f^{\prime}({\bm{x}}^{\prime},\cdot)}))$
@@
Hence, for the function $f({\bm{x}},{\bm{w}})$ with respect to ${\bm{w}}$, each element ${\bm{x}}\in\mathcal{X}_{t}\cup\{\hat{\bm{x}}_{t}\}$ satisfies ${f({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})$. Moreover, from the theorem’s assumption, since ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ satisfy (3.2), the following inequality holds:
@@
where the third inequality is derived by ${\mu_{t-1}({\bm{x}},\cdot)}\in G_{t-1}({\bm{x}})$, the definition of $\hat{\bm{x}}_{t}$ and $\max_{{\bm{x}}\in\mathcal{X}}{\rm lcb}_{t-1}({\bm{x}})\leq\max_{{\bm{x}}\in\mathcal{X}}\rho({\mu_{t-1}({\bm{x}},\cdot)})=\rho({\mu_{t-1}(\hat{\bm{x}}_{t},\cdot)})$. Furthermore, from Assumption 4.1, there exists a function
@@
$F({\bm{x}}^{\ast})-F([{\bm{x}}^{\ast}]_{t})=\rho({f({\bm{x}}^{\ast},\cdot)})-\rho({f([{\bm{x}}^{\ast}]_{t},\cdot)})\leq q_{1}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f([{\bm{x}}^{\ast}]_{t},{\bm{w}}))|\right).$
@@ C.7 Upper Bound of $\mathbb{E}[L_{max}]$
Thus, for the function ${f^{\dagger}_{t}({\bm{x}},{\bm{w}})}$ with respect to ${\bm{w}}$, ${f^{\dagger}_{t}({\bm{x}},\cdot)}\in G^{\dagger}_{t-1}({\bm{x}})$ holds. Here, from the theorem’s assumption, since ${\rm ucb}^{\dagger}_{t-1}({\bm{x}})$ and ${\rm lcb}^{\dagger}_{t-1}({\bm{x}})$ satisfy (A.1), the following holds:
@@
$F({\bm{x}}^{\ast})-F^{\dagger}_{t}({\bm{x}}^{\ast})=\rho({f({\bm{x}}^{\ast},\cdot)})-\rho({f^{\dagger}_{t}({\bm{x}}^{\ast},\cdot)})\leq q_{2}\left(\max_{{\bm{w}}\in\Omega}|f({\bm{x}}^{\ast},{\bm{w}}))-f({\bm{x}}^{\ast},[{\bm{w}}]_{t}))|\right).$
@@
$=\mathbb{E}_{{\bm{w}}}[\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},{\bm{w}})]=\rho({\mu^{\dagger}_{t-1}(\hat{\bm{x}}^{\dagger}_{i},\cdot)}).$
@@
By combining these, and using the same argument as in the proof of Theorem 4.1, we obtain Theorem B.1. Theorems B.2B.4 can also be obtained by using the same argument. ∎
C.12 Proof of Theorems B.5B.8
@@
To show Theorems B.5B.8, we evaluate $\mathbb{E}[\sigma^{2}_{k-1}({\bm{x}}_{k},{\bm{w}}^{(max)}_{k})]$ by using the same argument as in the proof of Theorems B.1B.4. Under the given
@@ D Experimental Details and Additional Experiments
In this section, we give details of the numerical experiments.
@@ D.1 Experimental Details
BPT-UCB (fixed)
In BPT-UCB (fixed), the definition of ${\rm BPTUCB}({\bm{x}})$ in BPT-UCB was changed to ${\rm BPTUCB}({\bm{x}})=\hat{p}_{t}({\bm{x}})+3\gamma_{t}({\bm{x}})$, and ${\bm{x}}_{t+1}$ and ${\bm{w}}_{t+1}$ were selected using the same procedure as BPT-UCB.
@@
BBBMOBO (fixed)
In BBBMOBO (fixed), we used $\beta_{t}=9$ and ${\bm{x}}_{t}=\tilde{\bm{x}}_{t}$, where $\tilde{\bm{x}}_{t}$ is given in Definition 3.1.
@@
Proposed (fixed)
In Proposed (fixed), ${\bm{x}}_{t}$ was selected by Definition 3.1, where we used $\beta_{t}=9$.
Values of ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ in EXP, PTR and EXP-MAE were calculated by using results in Table 3 in Inatsu et al., 2024a . Here, ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for $-\alpha\mathbb{E}[|f({\bm{x}},{\bm{w}})-F_{{\rm exp}}({\bm{x}})|]$ were calculated by using the results of the mean absolute deviation and monotonic Lipschitz map in the table. Combining these and the results for the weighted sum, we calculated ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ for EXP-MAE.
D.2 Identity of BBBMOBO (fixed) and Proposed (fixed) in the Expectation Measure
In the expectation measure, ${\rm ucb}_{t-1}({\bm{x}})$ and ${\rm lcb}_{t-1}({\bm{x}})$ are calculated as follows:
${\rm ucb}_{t-1}({\bm{x}})=\mathbb{E}_{{\bm{w}}}[\mu_{t-1}({\bm{x}},{\bm{w}})+\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},{\bm{w}})],\ {\rm lcb}_{t-1}({\bm{x}})=\mathbb{E}_{{\bm{w}}}[\mu_{t-1}({\bm{x}},{\bm{w}})-\beta^{1/2}_{t}\sigma_{t-1}({\bm{x}},{\bm{w}})].$
Therefore, (3.4) in Definition 3.1 is given by ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}}\mathbb{E}_{\bm{w}}[\sigma_{t-1}({\bm{x}},{\bm{w}})]$. Also, the definition of $\tilde{\bm{x}}_{t}$ can be rewritten as $\tilde{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}{\rm ucb}_{t-1}({\bm{x}})$. Here, the difference between BBBMOBO (fixed) and Proposed (fixed) is only whether to use ${\bm{x}}_{t}=\tilde{{\bm{x}}}_{t}$ or (3.4) in Definition 3.1. In Proposed (fixed), when ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}}\mathbb{E}_{\bm{w}}[\sigma_{t-1}({\bm{x}},{\bm{w}})]=\tilde{\bm{x}}_{t}$, it is consistent with BBBMOBO (fixed). Also, when ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}}\mathbb{E}_{\bm{w}}[\sigma_{t-1}({\bm{x}},{\bm{w}})]=\hat{\bm{x}}_{t}$,
${\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})$ $=\mathbb{E}_{\bm{w}}[\mu_{t-1}(\tilde{\bm{x}}_{t},{\bm{w}})+\beta^{1/2}_{t}\sigma_{t-1}(\tilde{\bm{x}}_{t},{\bm{w}})]=\mathbb{E}_{\bm{w}}[\mu_{t-1}(\tilde{\bm{x}}_{t},{\bm{w}})]+\beta^{1/2}_{t}\mathbb{E}_{\bm{w}}[\sigma_{t-1}(\tilde{\bm{x}}_{t},{\bm{w}})]$
$\leq\mathbb{E}_{\bm{w}}[\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}})]+\beta^{1/2}_{t}\mathbb{E}_{\bm{w}}[\sigma_{t-1}(\tilde{\bm{x}}_{t},{\bm{w}})]$
$\leq\mathbb{E}_{\bm{w}}[\mu_{t-1}(\hat{\bm{x}}_{t},{\bm{w}})]+\beta^{1/2}_{t}\mathbb{E}_{\bm{w}}[\sigma_{t-1}(\hat{\bm{x}}_{t},{\bm{w}})]={\rm ucb}_{t-1}(\hat{\bm{x}}_{t}).$
Here, the first inequality follows from the definition of $\hat{\bm{x}}_{t}$, $\hat{\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\mathcal{X}}\mathbb{E}_{\bm{w}}[\mu_{t-1}({\bm{x}},{\bm{w}})]$, and the second inequality is derived by the assumption, ${\bm{x}}_{t}=\operatornamewithlimits{argmax}_{{\bm{x}}\in\tilde{\bm{x}}_{t},\hat{\bm{x}}_{t}}\mathbb{E}_{\bm{w}}[\sigma_{t-1}({\bm{x}},{\bm{w}})]=\hat{\bm{x}}_{t}$. Therefore, ${\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})\leq{\rm ucb}_{t-1}(\hat{\bm{x}}_{t})$, but on the other hand, from the definition of $\tilde{\bm{x}}_{t}$, ${\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})\geq{\rm ucb}_{t-1}(\hat{\bm{x}}_{t})$. Hence, ${\rm ucb}_{t-1}(\tilde{\bm{x}}_{t})={\rm ucb}_{t-1}(\hat{\bm{x}}_{t})$. In the EXP experiments in Section 5, there were no cases where multiple ${\bm{x}}$ values maximized ${\rm ucb}_{t-1}({\bm{x}})$. As a result, $\tilde{\bm{x}}_{t}=\hat{\bm{x}}_{t}$, which is consistent with BBBMOBO (fixed).